# Accretion of a Vlasov gas by a Kerr black hole

### v1

@@ VI.1 Monoergetic models
In Figure 14 we show the dependence of $\dot{\mathcal{N}}$, computed according to Eq. (80), on the black hole spin parameter $\alpha$. The left panel shows the behavior characteristic for the energies $\varepsilon_{0}$ close to one. In this case, we plot $\dot{\mathcal{N}}\sqrt{\varepsilon_{0}^{2}-1}/(n_{\infty}M^{2})$, which tends to a finite limit as $\varepsilon_{0}\to 1$. The right panel illustrates the behavior of $\dot{\mathcal{N}}/(n_{\infty}M^{2})$ for large energies $\varepsilon_{0}$. In all cases, when $\alpha=0$, we recover the values known from spherical models corresponding to the Schwarzschild background spacetime.
@@ VI.2 Maxwell-Jüttner models
In all cases, $\dot{\mathcal{N}}/(n_{\infty}M^{2})$ and $\dot{\mathcal{E}}/(\varepsilon_{\infty}M^{2})$ decrease with $\alpha$. This remains in agreement with the behavior observed for razor-thin accretion disks confined to the equatorial plane and investigated in [25], or with the properties of accretion flows in the Reissner-Nordström spacetime, understood as a toy-model for the Kerr geometry [11]. The same behavior is also shared by hydrodynamical models [5]. In contrast to planar models of [25], both quantities $\dot{\mathcal{N}}/(n_{\infty}M^{2})$ and $\dot{\mathcal{E}}/(\varepsilon_{\infty}M^{2})$ decrease with increasing asymptotic temperature (they increase with $z$). This is again to be expected, as this behavior agrees with the properties of spherically symmetric flows in the Schwarzschild spacetime [9, 10].
The curves presented in Figs. 1417 correspond to the approximated particle and energy accretion rates given by Eqs. (91) and (92) for monoenergetic and Maxwell-Jüttner models. The difference between the results obtained within the slow-rotation approximation and the exact numerical values is smaller than $1\%$ for rotating Kerr black holes with $\alpha\leq 0.8$ and remains smaller than $4\%$ even for the rapidly rotating case with $\alpha\leq 0.99$. [^7] For better visibility, some curves corresponding to the slow-rotation approximation are omitted in Figs. 1416. We have checked that the statement regarding the difference between the slow-rotation approximation and the exact numerical values applies to the missing curves as well. Thus, in practical applications it should be possible to replace the original triple integrals appearing in Eqs. (81) with simpler formulas (92) obtained in the slow-rotation approximation, without introducing a significant error.
@@ VII Conclusions
In this work, we obtained new solutions of the Vlasov equation describing the Bondi-type accretion of a collisionless gas onto a rotating Kerr black hole. Deriving such solutions requires good control of the regions in the phase space available for the motion of the gas particles. Results presented in this work are based on a characterization of these regions in terms of specially adapted “momentum” coordinates $E$, $Q$, and $\chi$. [^8] When this work was nearly completed, Ref. [50] appeared which treats the Bondi-type accretion problem in a Kerr-Newman black hole and uses a parametrization which, in the Kerr limit, reduces to the same variables $Q$ and $\chi$ used in this work. However, the author in [50] seems to reach different conclusions regarding the accretion rates; in particular, a non-vanishing angular momentum accretion rate is obtained even in the Kerr case.
@@
As particular examples, we considered two models, differing in the assumed asymptotic form of the one-particle distribution function: a monoenergetic distribution and a Maxwell-Jüttner distribution function characterized by constant temperature and constant particle density at infinity. In both cases, the particle and energy accretion rates decrease with the growing black hole spin parameter. Quite surprisingly, the angular momentum accretion accretion rate vanishes for all models. This stays in contrast with the model of thin accretion disks confined to the equatorial plane, investigated in [25, 26], and can be explained by cancellations occurring in the momentum space. In the quasistationary approximation (i.e., with the evolution of the system approximated by a sequence of stationary solutions corresponding to different black hole parameters), such behavior should lead to a decrease in the black hole spin parameter $\alpha$, due to an increase of the black hole mass alone.
@@
Particle and energy accretion rates can also be computed assuming a slow-rotation approximation. Many of the results obtained within the slow-rotation approximation agree with the exact solution, even for large values of the black hole spin parameters. This applies especially to particle and energy accretion rates, but also for components of the particle current density, except for a region close to the black hole horizon.
Many natural problems associated with the model presented in this paper are left for future studies. They include addressing the questions about the mass inflation at the Cauchy horizon or taking into account the effects of self-gravity of the gas. Another natural generalization involves considering a uniform velocity of the gas at infinity (accretion onto a moving black hole), especially in the case in which the asymptotic velocity of the gas is misaligned with the rotation axis. That, in turn, implies natural questions about drag effects (see, e.g., [51, 52]).
@@ Acknowledgements.
P. M. acknowledges a support of the Polish National Science Centre Grant No. 2017/26/A/ST2/00530. M. M. was supported by CONAHCyT through Estancias Posdoctorales por México Convocatoria 2023(1) under the postdoctoral Grant No. 1242413. O. S. was partially supported by CIC Grant No. 18315 to Universidad Michoacana and by CONAHCyT Network Project No. 376127 “Sombras, lentes y ondas gravitatorias generadas por objetos compactos astrofísicos”. Finally, M. M. and O. S. acknowledge financial support from SECIHTI-SNII.
@@ Appendix D On the smoothness of the observables
for $r$ lying close to $r_{c}$, where in the second step we have use the fact that the integrand is an even function of $\eta$. Note that both $c_{1}$ and $c_{2}$ remain finite. Moreover, at $r=r_{c}$, we have $A(Q_{+}-Q_{-})=A/c_{2}=\kappa^{2}$. Since the right-hand side of Eq. (161) has the form of a sign function multiplied with a smooth integral, it follows that its first derivative must jump at $r=r_{c}$.

### v3

@@ VI.1 Monoergetic models
Figure 15: Angular momentum accretion rates for monoenergetic models. We plot $\dot{\mathcal{J}}\sqrt{\varepsilon_{0}^{2}-1}/(\varepsilon_{\infty}M^{3})$ for the energies $\varepsilon_{0}$ close to $\varepsilon_{0}=1$ (left panel) and $\dot{\mathcal{J}}/(\varepsilon_{\infty}M^{3})$ for large energies (right panel). For a rapidly rotating Kerr black hole with $\alpha\leq 0.99$, the error is less than $5\%$ in the left panel and less than $2.7\%$ in the right panel.
@@
In Figs. 1415, we show the dependence of $\dot{\mathcal{N}}$ and $\dot{\mathcal{J}}$, computed according to Eq. (80), on the black hole spin parameter $\alpha$. The left panels show the behavior characteristic for the energies $\varepsilon_{0}$ close to one. In this case, we plot $\dot{\mathcal{N}}\sqrt{\varepsilon_{0}^{2}-1}/(n_{\infty}M^{2})$ and $\dot{\mathcal{J}}\sqrt{\varepsilon_{0}^{2}-1}/(\varepsilon_{\infty}M^{3})$, which tends to a finite limit as $\varepsilon_{0}\to 1$. The right panels illustrate the behavior of $\dot{\mathcal{N}}/(n_{\infty}M^{2})$ and $\dot{\mathcal{J}}/(\varepsilon_{\infty}M^{3})$ for large energies $\varepsilon_{0}$. In all cases, when $\alpha=0$, we recover the values known from spherical models corresponding to the Schwarzschild background spacetime.
@@ VI.2 Maxwell-Jüttner models
Figure 19: The angular momentum accretion rates for the Maxwell-Jüttner distribution and various values of $z$ as a function of $\alpha$. For all rotating Kerr black holes with $\alpha\leq 0.99$, the error is less than $2.6\%$ in the left panel and less than $5\%$ in the right panel.
Figure 20: Compression ratio $n/n_{\infty}$ versus radius $x$ in the low-temperature limit at the equatorial plane ($\vartheta=\pi/2$). Discrete marks (plotted for $x>1$) correspond to exact numerical values obtained for $z=200$. The analytic relation (95) of the slow-rotation approximation is plotted with lines for $x\geq 1.3$. Vertical dashed lines mark locations of the event horizon $x_{+}$. At $x=1.5$, the difference between approximated and exact values is less than $1.5\%$ (inside the horizon) for $\alpha=0.5$ and smaller than $12\%$ (outside the horizon) for $\alpha=0.95$.
@@
In all cases, $\dot{\mathcal{N}}/(n_{\infty}M^{2})$ and $\dot{\mathcal{E}}/(\varepsilon_{\infty}M^{2})$ decrease with $\alpha$, whereas $\dot{\mathcal{J}}/(\varepsilon_{\infty}M^{3})$ increases. This remains in agreement with the behavior observed for razor-thin accretion disks confined to the equatorial plane and investigated in [27], or with the properties of accretion flows in the Reissner-Nordström spacetime, understood as a toy-model for the Kerr geometry [11]. The same behavior is also shared by hydrodynamical models [5]. In contrast to planar models of [27], both quantities $\dot{\mathcal{N}}/(n_{\infty}M^{2})$ and $\dot{\mathcal{E}}/(\varepsilon_{\infty}M^{2})$ decrease with increasing asymptotic temperature (they increase with $z$). This is again to be expected, as this behavior agrees with the properties of spherically symmetric flows in the Schwarzschild spacetime [9, 10]. However, $\dot{\mathcal{J}}/(\varepsilon_{\infty}M^{3})$ is a decreasing function of $z$.
The curves presented in Figs. 1419 correspond to the approximated particle, energy, and angular momentum accretion rates given by Eqs. (91) and (92) for monoenergetic and Maxwell-Jüttner models. The difference between the results obtained within the slow-rotation approximation and the exact numerical values is smaller than $1\%$ ($1.1\%$ in case of $\mathcal{\dot{J}}$) for rotating Kerr black holes with $\alpha\leq 0.8$ and remains smaller than $4\%$ ($5\%$ in case of $\mathcal{\dot{J}}$) even for the rapidly rotating case with $\alpha\leq 0.99$. [^7] For better visibility, some curves corresponding to the slow-rotation approximation are omitted in Figs. 1417 and 19. We have checked that the statement regarding the difference between the slow-rotation approximation and the exact numerical values applies to the missing curves as well. Thus, in practical applications it should be possible to replace the original triple integrals appearing in Eqs. (81) with simpler formulas (92) obtained in the slow-rotation approximation, without introducing a significant error.
Finally, in Fig. 20, we plotted the compression ratio $n/n_{\infty}$ versus radius $x$ in the low-temperature limit given by the analytic formula (95) along with the corresponding exact numerical values. This figure shows that for points in the equatorial plane which have $x\geq 1.5$, the difference in $n/n_{\infty}$ between the results obtained within the slow-rotation approximation and the exact numerical values is smaller than $1.5\%$ for $\alpha=0.5$ and smaller than $12\%$ for $\alpha=0.95$.
@@ VII Conclusions
In this work, we obtained new solutions of the Vlasov equation describing the Bondi-type accretion of a collisionless gas onto a rotating Kerr black hole. Deriving such solutions requires good control of the regions in the phase space available for the motion of the gas particles. Results presented in this work are based on a characterization of these regions in terms of specially adapted “momentum” coordinates $E$, $Q$, and $\chi$. [^8] When this work was nearly completed, Ref. [53] appeared which treats the Bondi-type accretion problem in a Kerr-Newman black hole and uses a parametrization which, in the Kerr limit, reduces to the same variables $Q$ and $\chi$ used in this work. However, the author in [53] reaches different conclusions regarding the accretion rates.
@@
As particular examples, we considered two models, differing in the assumed asymptotic form of the one-particle distribution function: a monoenergetic distribution and a Maxwell-Jüttner distribution function characterized by constant temperature and constant particle density at infinity. In both cases, the particle and energy accretion rates decrease with the growing black hole spin parameter. We also find that the angular momentum accretion rate is roughly proportional to the spin parameter, slowing the black hole rotation down. This stays in agreement with the model of thin accretion disks confined to the equatorial plane, investigated in [27, 28]. In the quasistationary approximation (i.e., with the evolution of the system approximated by a sequence of stationary solutions corresponding to different black hole parameters), this effect, together with an increase of the black hole mass, should lead to a decrease in the black hole spin parameter $\alpha$.
@@
Particle, energy, and the angular momentum accretion rates can also be computed assuming a slow-rotation approximation. Many of the results obtained within the slow-rotation approximation agree well with the exact solution, even for large values of the black hole spin parameters. This applies not only to the accretion rates, but also for components of the particle current density, except for a region close to the black hole horizon.
Many natural problems associated with the model presented in this paper are left for future studies. They include addressing questions about mass inflation at the Cauchy horizon or taking into account the effects of self-gravity of the gas. Taking into account the self-gravity (i.e., solving the full Einstein–Vlasov problem) requires addressing several issues. On one hand, different types of particle orbits can no longer be analyzed independently. In particular an inclusion of particles on bound orbits would affect particles associated with unbound ones. On the other hand, in a Bondi-type model a self-gravitating gas cannot extend to infinity. In spherical symmetry, hydrodynamical Bondi-type accretion of a self-gravitating fluid was investigated in [54, 55, 56, 57], assuming boundary conditions imposed at a finite radius. Imposing boundary conditions at a finite boundary introduces a modification even in kinetic models with a fixed spacetime background (see, e.g., [17, 28]).
Since in this work we consider an unmagnetized gas of neutral particles our results cannot be applied to astrophysical plasmas. However, they do apply to the accretion of dark matter or distributions of stars. Some examples of dark matter models based on the kinetic approach can be found in [58, 59]. We discuss two short applications to the accretion of dark matter on black holes in the accompanying paper [60], both for hot (in the early universe) and cold dark matter. In this regard, another natural generalization of our model involves considering a uniform velocity of the gas at infinity (accretion onto a moving black hole), especially in the case in which the asymptotic velocity of the gas is misaligned with the rotation axis. That, in turn, implies natural questions about drag effects (see, e.g., [61, 62]).
@@ Acknowledgements.
P. M. acknowledges a support of the Polish National Science Centre Grant No. 2017/26/A/ST2/00530. M. M. was supported by SECIHTI through Estancias Posdoctorales por México Convocatoria 2023(1) under the postdoctoral Grant No. 1242413. O. S. was partially supported by CIC Grant No. 18315 to Universidad Michoacana and by CONAHCyT Network Project No. 376127 “Sombras, lentes y ondas gravitatorias generadas por objetos compactos astrofísicos”. Finally, M. M. and O. S. acknowledge financial support from SECIHTI-SNII. The data that support the findings of this article are openly available [63].
@@ Appendix D On the smoothness of the observables
for $r$ lying close to $r_{c}$, where in the second step we have used the fact that the integrand is an even function of $\eta$. Note that both $c_{1}$ and $c_{2}$ remain finite. Moreover, at $r=r_{c}$, we have $A(Q_{+}-Q_{-})=A/c_{2}=\kappa^{2}$. Since the right-hand side of Eq. (160) has the form of a sign function multiplied with a smooth integral, it follows that its first derivative must jump at $r=r_{c}$.